### v1

# Asymptotic Consistency and Generalization in Hybrid Models of Regularized Selection and Nonlinear Learning

Abstract
This study explores how different types of supervised models perform in the task of predicting and selecting relevant variables in high-dimensional contexts, especially when the data is very noisy. We analyzed three approaches: regularized models (such as Lasso, Ridge, and Elastic Net), black-box models (such as Random Forest, XGBoost, LightGBM, CatBoost, and H2O GBM), and hybrid models that combine both approaches: regularization with nonlinear algorithms. Based on simulations inspired by the Friedman equation, we evaluated 23 models using three complementary metrics: RMSE, Jaccard index, and recall rate. The results reveal that, although black-box models excel in predictive accuracy, they lack interpretability and simplicity, essential factors in many real-world contexts. Regularized models, on the other hand, proved to be more sensitive to an excess of irrelevant variables. In this scenario, hybrid models stood out for their balance: they maintain good predictive performance, identify relevant variables more consistently, and offer greater robustness, especially as the sample size increases. Therefore, we recommend using this hybrid framework in market applications, where it is essential that the results make sense in a practical context and support decisions with confidence.
Key Words: Regularization; Black-box models; Variable selection; High dimensionality; RMSE; Jaccard; Model interpretation.
1  Introduction
Predictive models for high-dimensional problems pose a significant computational challenge. In many cases, the direct use of nonlinear algorithms to reduce the number of features can be computationally expensive. Although widely used, this framework can generate biases, overfitting (induced by heuristics), and limitations in methods based on variable importance, given that the predictive model itself influences the prioritization of selected variables [1, 2]. The use of non-heuristic methods, such as regularizations independent of the predictive model, has gained prominence in several fields to mitigate such biases, with hybrid models being a promising solution [3]. However, these associated models require further theoretical development regarding their generalization capacity and asymptotic properties.
This work attempts to develop a theoretical framework that analyzes the selection consistency and generalization limits of these hybrid models, focusing on the combination of regularized methods such as Lasso, Ridge, and Elastic Net with widely used black-box algorithms, including Random Forest, XGBoost, LightGBM, CatBoost, and H2O GBM.
Regarding the selection step using regularized models, Lasso ([4, 5]) is a selection and regularization method that promotes sparsity in the coefficients through the penalty $\ell 1$. Lasso’s consistency in recovering the true set of variables depends on conditions such as unrepresentability and strong signal strength, which are essential in high-dimensional scenarios. Ridge Regression ([6]) applies quadratic penalty ($\ell 2$) and does not perform variable selection, but contributes to greater stability and variance reduction in the presence of multicollinearity. Elastic Net [7] combines the penalties $\ell 1$ and $\ell 2$, offering a robust alternative to Lasso, especially effective in the face of high correlation between predictors.
Black-box models are widely used in supervised learning applications due to their high predictive capacity and flexibility. Random Forest is based on bagging [8], a technique that reduces variance by combining decision trees generated on bootstrapped subsets. Boosting-based models, such as XGBoost, LightGBM, and CatBoost, use sequential additive learning to reduce bias and improve performance, while presenting challenges in terms of stability and overfitting [9]. H2O GBM offers a scalable implementation of boosting with automatic hyperparameter adjustments. The theory underlying these models includes interpretations such as adaptive neighbors for Random Forest [10] and bias-variance analysis for more randomized ensembles such as extra-trees [11].
Algorithmic stability plays a central role in theoretical analysis and is directly connected to the behavior of generalization errors, as discussed by Bousquet and Elisseeff. In this context, the notion of expected risk consistency stands out, related to the Empirical Risk Minimization Principle (ERM), as presented by Vapnik. This principle establishes that a good model must present asymptotic convergence between the empirical risk observed in the data and the expected risk associated with the true function that generated the data. Such convergence is essential in supervised learning scenarios, as it ensures that, as the number of observations increases, the model’s performance tends to approach optimality within the class of functions considered.
This property is particularly relevant in scenarios where predictive accuracy is more important than the exact recovery of relevant variables, but here it is used to demonstrate the assertive power of the proposed hybrid models. In this context, even without identifying all the true support, a model can be considered predictive consistent if its predictive ability remains robust as the number of observations increases. Formal discussions of this form of consistency can be found in Vapnik, Bousquet and Elisseeff, and Bousquet and Elisseeff. The loss function and the type of penalty also impact the bias and variance of the estimator, Rosasco2004.
To empirically evaluate the proposed framework, controlled simulations were performed based on the Friedman equation, Friedman1991. This equation is widely used in the literature to generate synthetic data with a known ground truth structure, including relevant predictor variables and noise components. The experimental design allows a clear distinction between the true model and the estimated model, enabling the evaluation of method performance in different dimensionality regimes (n and p) and under different levels of noise and structural complexity of the data. Despite the vast literature on regularized methods and black-box models individually, there is a lack of a theoretical framework that rigorously connects and grounds these steps in the hybrid context. This article aims to contribute to filling this gap by offering new theoretical constraints and conditions for the consistency and generalization of the proposed hybrid pipeline.
2  Methodology
Variable selection in high-dimensional scenarios requires formal criteria that guarantee the theoretical validity of the methods used, especially when seeking to ensure stable performance in large samples. In this section, we discuss two fundamental concepts in this context: selection consistency, which ensures the recovery of true support, and expected risk consistency, which ensures the convergence of the prediction error to the optimal risk of the generating model.
These properties are essential for the asymptotic robustness of methods that combine regularization and nonlinear modeling, such as the hybrid frameworks proposed in this study. The detailed structure of the experiment is presented in Algorithm 1.
Algorithm 1 Simulation and Evaluation Framework with Regularization and Hybrid Models
1:Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathbb{R}$; sample sizes $n\in\{200,500,1000\}$; number of predictors $p\in\{5,10,50\}$; replicates $n_{\text{sim}}$; performance metrics: RMSE, Jaccard index, Recovery.
2:for sim = 1 to $n_{\text{sim}}$ do
3:  Generate synthetic data:
$x_{ij}\sim\mathcal{U}(0,1),\quad y_{i}=10\sin(\pi x_{i1}x_{i2})+20(x_{i3}-0.5)^{2}+10x_{i4}+5x_{i5}+\varepsilon_{i},\quad\varepsilon_{i}\sim\mathcal{N}(0,1)$
4:  for each model in {Ridge, Lasso, Elastic Net} do
5:    Set $\alpha=0$ (Ridge), $\alpha=1$ (Lasso), $0<\alpha<1$ (Elastic Net).
6:    Fit cv.glmnet with family = "gaussian":
  • Use MSE to select $\lambda_{\min}$.
  • Extract coefficients $\hat{\beta}_{j}$ at $\lambda_{\min}$.
7:    Rank variables by $|\hat{\beta}_{j}|$, forming $\hat{S}_{\text{ranked}}$.
8:    for m = 1 to min(10, p - 1) do
9:     Fit OLS on top-m variables from $\hat{S}_{\text{ranked}}$.
10:     Evaluate RMSE using k-fold CV.
11:    end for
12:    Select $m^{*}$ with minimal RMSE and define $\hat{S}=\hat{S}_{\text{ranked}}[1:m^{*}]$.
13:    Compute:
$\text{RMSE}_{\text{reg}},\quad\text{Jaccard}=\frac{|\hat{S}\cap S|}{|\hat{S}\cup S|},\quad\text{Recovery}=\frac{|\hat{S}\cap S|}{|S|}$
14:  end for
15:  for each black-box model in {RF, XGBoost, LightGBM, CatBoost, H2O GBM} do
16:    Fit model using all predictors to extract feature importance.
17:    Rank variables by importance and form $\hat{S}_{\text{imp}}$ with up to 10 variables.
18:    for m = 1 to min(10, p - 1) do
19:     Evaluate RMSE via k-fold CV on top-m variables from $\hat{S}_{\text{imp}}$.
20:    end for
21:    Select $m^{*}$ with minimal RMSE and define $\hat{S}_{\text{bb}}=\hat{S}_{\text{imp}}[1:m^{*}]$.
22:    Compute $\text{RMSE}_{\text{bb}}$, Jaccard and Recovery using $\hat{S}_{\text{bb}}$.
23:  end for
24:  for each combination of black-box and regularized subset do
25:    Apply black-box model on $\hat{S}$ from regularized model (hybrid).
26:    Compute RMSE, Jaccard and Recovery for hybrid.
27:  end for
28:end for
29:Return all stored metrics (RMSE, Jaccard, Recovery) across simulations.
Note. We set the upper bound of $m$ to $\min(10,p-1)$ to avoid trivial subsets including all predictors, which would artificially inflate support recovery metrics.
Next, we formalize some of the main definitions in the general context of the regression problem according to the proposed framework.
Definition: Asymptotic Consistency
Let $S\subset\{1,\dots,p\}$ be the true set of relevant variables and $\hat{S}$ be the set estimated by the model. We say that a variable selection method is selection consistent if
$P\left(\hat{S}=S\right)\xrightarrow[n\to\infty]{}=1.$ (1)
Furthermore, we say that the predictive estimator $\hat{f}$ is predictive consistent if
$\mathbb{E}\left[\left(f(\mathbf{X})-\hat{f}(\mathbf{X})\right)^{2}\right]\xrightarrow[n\to\infty]{}=0,$ (2)
where $f(\mathbf{X})$ is the truth function of the generative model.
Definition: Generalization
Let R(f^)R($\hat${f}) be the expected risk in the population:
$R(\hat{f})=\mathbb{E}{(Y,\mathbf{X})}\left[L(Y,\hat{f}(\mathbf{X}))\right],$ (3)
where $L(\cdot)$ is an appropriate loss function (e.g., squared error, deviance, log-loss). The model generalizes well if
$R(\hat{f})-R(f^{*})\xrightarrow[n\to\infty]{}0,$ (4)
where $f^{*}$ is the minimizer of the expected risk.
Consistency Proposition of the Hybrid Framework.
Suppose that the true model has a sparse structure with support $S\subset\{1,\dots,p\}$, and that the regularized selection step satisfies the conditions of unrepresentability and signal strength, ensuring $\mathbb{P}(\hat{S}=S)\to 1$ when $n\to\infty$. Further suppose that the predictive model fitted over the selected subspace (black-box model) is consistently predictive in the sense of $R(\hat{f})\to R(f^{*})$. Under these assumptions, the resulting hybrid framework tends to be consistent in both support recovery and predictive risk.
$\mathbb{P}(\hat{S}=S)\to 1\quad\text{e}\quad R(\hat{f})\to R(f^{*})$ (5)
This formulation does not constitute a formal theorem in the literature, but expresses a principle proposed in this work based on two widely recognized theoretical pillars: the consistency of selection via regularization and the predictive consistency of models fitted on relevant subsets [12].
2.1 Experimental Study
To empirically validate our findings, we chose to conduct simulation studies. The simulations were based on the Friedman model, widely used in the literature to evaluate variable selection methods in nonlinear and high-dimensional scenarios. What is considered a true model corresponds to the case in which the predictor vector $\mathbf{x}_{i}\in\mathbb{R}^{p}$ has real support $S=\{1,2,3,4,5\}$, meaning that only the first five variables affect the response.
In this context, the model with additive noise is defined by:
$Y_{i}=f(\mathbf{x}_{i})+\varepsilon_{i}=10\sin(\pi x_{i1}x_{i2})+20(x_{i3}-0.5)^{2}+10x_{i4}+5x_{i5}+\varepsilon_{i},$ (6)
where $\varepsilon_{i}\sim\mathcal{N}(0,1)$ represents the error term, and each covariate $x_{ij}\overset{\text{i.i.d.}}{\sim}\mathcal{U}(0,1),\ for\ j=1,\dots,p$.
Although only the first five variables are present in the response generating function $f(\mathbf{x}_{i})$, the vector $\mathbf{x}_{i}$ contains up to $p=50$ covariates, of which p5p-5 are irrelevant and represent stochastic noise in the model.
To evaluate the methods’ ability to recover the true set of relevant variables, we also consider a deterministic version of the model, that is, without additive noise, defined by:
$Y_{i}=f(\mathbf{x}_{i})=10\sin(\pi x_{i1}x_{i2})+20(x_{i3}-0.5)^{2}+10x_{i4}+5x_{i5}.$ (7)
This noise-free model allows for the isolated measurement of the exact recovery capacity of variable selection methods, as it eliminates the interference of random error.
2.2 Models Evaluated
In this study, three main groups of models were considered. The first group comprises the regularized models: Ridge (with penalty $\alpha=0$), Lasso ($\alpha=1$), and Elastic Net ($\alpha=0.5$), fitted using the glmnet package [13]. Variable selection in these models was performed based on the support of the non-zero coefficients obtained at the penalty point $\lambda_{\min}$, chosen via cross-validation.
The second group includes black-box models, which include Random Forest [14], XGBoost [15], LightGBM [16], CatBoost [17], and H2O GBM. These algorithms were fitted on all available predictors, without a prior selection step.
Finally, the third group consists of hybrid models, in which each black-box model was fitted exclusively on the subset of variables previously selected by one of the regularized models, seeking to combine parsimony with predictive flexibility.
2.3 Simulations
This study evaluated the performance of regularized, pure, and hybrid black-box models under twelve distinct scenarios, defined by combinations of three sample sizes ($n\in\{200,500,1000\}$) and four dimensionality levels ($p\in\{5,10,50,100\}$). For each scenario, ten simulated datasets were generated independently, totaling 120 replications in total.
For each replication, 23 models were applied: 3 regularized models (Ridge, Lasso, and Elastic Net), 5 pure black-box models (Random Forest, XGBoost, LightGBM, CatBoost, and H2O GBM), and 15 hybrid models resulting from the combination of regularizers and black-box algorithms.
Thus, each scenario involved 230 runs (10 datasets × 23 models), totaling 2,760 runs throughout the simulation. Table 1 details the scenarios considered and the composition of the models evaluated by type.
Table 1: Simulation parameters, number of executions per scenario, and composition of model types.
$n$ $p$ Scenario Model Quantity
200 | 500 | 1000 5 230 Regularized 3
Pure black-box 5
Hybrids (Reg × BB) 15
200 | 500 | 1000 10 230 Regularized 3
Pure black-box 5
Hybrids (Reg × BB) 15
200 | 500 | 1000 50 230 Regularized 3
Pure black-box 5
Hybrids (Reg × BB) 15
200 | 500 | 1000 100 230 Regularized 3
Pure black-box 5
Hybrids (Reg × BB) 15
For each combination of sample size $n$ and number of predictors $p$, 10 independent simulations were conducted to ensure robustness in the evaluations. In each of these replications, the data were initially randomly partitioned into training (80%) and test (20%) sets.
The regularized models were then adjusted using the penalty corresponding to the optimal value $\lambda_{\text{min}}$, obtained by cross-validation using the lowest RMSE criterion per feature subset. Based on the set of variables selected by these models, the hybrid models were then adjusted, limiting themselves exclusively to the estimated support.
For comparison purposes, the pure black-box models were also adjusted, considering all available predictors, regardless of relevance. The final performance of the models was then evaluated based on the RMSE (Root Mean Squared Error), Jaccard index, and Recovery metrics.
To ensure comparability between the methods, the black-box algorithms were trained with 100 trees or iterations. In particular, the models based on the H2O package [18] were initialized with full parallelism to optimize computational performance.
In Section  3, we focus on the $n=\{200,500,1000\}$ values of the main results for summary and clarity, although the full range evaluated is documented in the appendices and supplementary scripts.
2.4 Evaluation Metrics
Model performance was evaluated using three metrics. The first is RMSE (Root Mean Squared Error), which quantifies the mean squared error of predictions on the test set and is widely used as a measure of predictive accuracy.
The other two metrics refer to the quality of variable selection: the Jaccard index [19], defined as
$J=\frac{|S\cap\hat{S}|}{|S\cup\hat{S}|}$,
and the recovery rate [20], given by
$R=\frac{|S\cap\hat{S}|}{|S|},$
where $S$ represents the true set of relevant variables and $\hat{S}$ the set estimated by the model. Both are widely adopted in the literature to assess the consistency of variable selection.
2.5 Computer Implementation
The simulations were implemented in the R [21] software, using the following packages: glmnet (regularized models) [22], randomForest (Random Forest) [23], xgboost [24], lightgbm [25], catboost [26], h2o (with parallelism enabled for H2O GBM) [18], dplyr [27], tibble [28], tidyr[29], Metrics[30] purrr[31], doParallel[32], and foreach (for parallelization of simulations).
The results were saved in .csv and .RData files for reproducibility. Graphical visualizations were generated with the ggplot2 and viridis packages.
3  Results
In this section, we will present the results obtained in the simulation studies. The evaluation compared different modeling strategies, including regularized approaches, black-box models, and hybrid models that combine selection via regularization with nonlinear predictors. The evaluation metrics considered were the root mean square error (RMSE), the Jaccard index (for similarity between selected variables and the true set), and the recovery rate of relevant variables. The models in the subsections are analyzed segmented, allowing us to evaluate the relative performance of each class of approach under different complexity and generalization criteria.
3.1 Pure Black-Box Models
Table 2 presents the results of the main black-box models in their pure form, denoted as Full because they use the same algorithm for both variable selection and prediction, without combining them with regularized methods. Five algorithms are considered: CatBoost, H2O GBM, LightGBM, Random Forest, and XGBoost. The comparison includes three metrics at different sample sizes: root mean square error (RMSE), Jaccard index, and recall rate.
Table 2: Best performing pure black-box models with 95% noise ($p=50$).
Model Metric $n=200$ $n=500$ $n=1000$
CatBoost_Full RMSE 2.02 (0.22) 1.68 (0.26) 1.51 (0.29)
CatBoost_Full Jaccard 0.85 (0.13) 0.89 (0.11) 0.93 (0.10)
CatBoost_Full Recovery 0.94 (0.09) 0.95 (0.09) 0.95 (0.09)
H2OGBM_Full RMSE 2.13 (0.20) 1.90 (0.19) 1.72 (0.21)
H2OGBM_Full Jaccard 0.88 (0.13) 0.83 (0.14) 0.85 (0.14)
H2OGBM_Full Recovery 0.95 (0.09) 0.95 (0.09) 0.95 (0.09)
LightGBM_Full RMSE 2.16 (0.19) 1.86 (0.21) 1.68 (0.23)
LightGBM_Full Jaccard 0.78 (0.16) 0.80 (0.13) 0.89 (0.12)
LightGBM_Full Recovery 0.95 (0.09) 0.95 (0.09) 0.95 (0.09)
RF_Full RMSE 2.52 (0.13) 2.21 (0.13) 1.99 (0.14)
RF_Full Jaccard 0.80 (0.07) 0.82 (0.01) 0.82 (0.01)
RF_Full Recovery 0.94 (0.10) 0.95 (0.09) 0.95 (0.09)
XGBoost_Full RMSE 2.33 (0.20) 2.03 (0.19) 1.81 (0.23)
XGBoost_Full Jaccard 0.86 (0.12) 0.88 (0.11) 0.92 (0.10)
XGBOOST _FULL RECOVERY  TEXTBF 0.94 (0.09)  TextBF 0.95 (0.09)  TextBF 0.95 (0.09)
The results demonstrated that black-box models stand out for their high predictive power, with RMSE values consistently lower than those of the other approaches. The CatBoost_Full model obtained the lowest errors across all sample sizes, with $RMSE=1.51$ when $n=1000$ standing out. It also presented high Jaccard and Recovery values, indicating stability in identifying relevant variables. In terms of overlap with the true support, the XGBoost_Full, H2OGBM_Full, and LightGBM_Full models also demonstrated excellent performance, with recovery rates equal to or close to $0.95$ in all scenarios. The Jaccard index, in turn, was also high for H2OGBM_Full ($J=0.88$ at $n=200$) and for XGBoost_Full ($J=0.92$ at $n=1000$).
Although effective in predictive terms, pure black-box models do not promote, a priori, parsimony and offer little transparency in variable selection. This represents a significant obstacle in contexts that require interpretation or validation of the predictors used.
3.2 Pure Regularized Models
Table 3 presents the performance of penalized models fitted directly to simulated data with high spurious dimensionality (p = 50). These include Ridge, Lasso, and Elastic Net, methods that seek to reconcile prediction and selection through coefficient regularization.
Table 3: Pure regularized models with 95% noise ($p=50$).
Model Metric $n=200$ $n=500$ $n=1000$
Ridge RMSE 2.64 (0.19) 2.48 (0.18) 2.24 (0.17)
Ridge Jaccard 0.61 (0.18) 0.66 (0.16) 0.65 (0.17)
Ridge Recovery 0.83 (0.09) 0.87 (0.07) 0.87 (0.07)
Lasso RMSE 2.73 (0.18) 2.51 (0.18) 2.25 (0.17)
Lasso Jaccard 0.63 (0.19) 0.68 (0.16) 0.66 (0.18)
Lasso Recovery 0.84 (0.08) 0.87 (0.07) 0.87 (0.07)
ElasticNet RMSE 2.67 (0.19) 2.50 (0.18) 2.26 (0.17)
ElasticNet Jaccard 0.62 (0.19) 0.67 (0.16) 0.66 (0.17)
ElasticNet Recovery 0.84 (0.08) 0.87 (0.07) 0.87 (0.07)
It can be observed that, even with an increase in the sample size, the regularized models show limitations in predictive performance, with RMSE values consistently higher than those of the hybrid and black-box models. Nevertheless, there is a clear trend of improvement with increasing $n$, indicating that these methods benefit from greater sample information, albeit less significantly. Regarding variable selection, the recovery rate remains robust and stable ($R\approx 0.87$), revealing a good ability to identify truly relevant variables.
However, the Jaccard index remains moderate ($J\approx 0.66$), suggesting that the models still select a considerable number of irrelevant variables, which may compromise parsimony and increase the risk of overfitting, especially in contexts with high noise.
3.3 Hybrid Models - regularized black boxes
Table  4 presents the highlighted results of the hybrid models, formed by an initial selection step using regularized regression (Lasso, Ridge or Elastic Net), followed by the fitting of a non-linear model to the selected subset of variables. In section 5, table LABEL:tab:hybrid_models_full presents the results in their entirety.
Table 4: Hybrid models (regularization + black-box) with $p=50$.
Model Metric $n=200$ $n=500$ $n=1000$
CatBoost_ElasticNet RMSE 2.40 (0.25) 2.11 (0.26) 1.93 (0.35)
CatBoost_ElasticNet Jaccard 0.61 (0.19) 0.61 (0.20) 0.61 (0.20)
CatBoost_ElasticNet Recovery 0.83 (0.07) 0.84 (0.08) 0.86 (0.09)
CatBoost_Lasso RMSE 2.37 (0.23) 2.15 (0.28) 1.97 (0.32)
CatBoost_Lasso Jaccard 0.62 (0.20) 0.61 (0.21) 0.61 (0.18)
CatBoost_Lasso Recovery 0.84 (0.08) 0.84 (0.08) 0.85 (0.08)
CatBoost_Ridge RMSE 2.40 (0.21) 2.15 (0.24) 1.89 (0.35)
CatBoost_Ridge Jaccard 0.61 (0.19) 0.59 (0.19) 0.62 (0.21)
CatBoost_Ridge Recovery 0.83 (0.07) 0.83 (0.07) 0.86 (0.09)
H2OGBM_Lasso RMSE 2.40 (0.21) 2.19 (0.18) 2.01 (0.22)
H2OGBM_Lasso Jaccard 0.62 (0.20) 0.61 (0.21) 0.61 (0.18)
H2OGBM_Lasso Recovery 0.84 (0.08) 0.84 (0.08) 0.85 (0.08)
H2OGBM_ElasticNet RMSE 2.42 (0.20) 2.18 (0.20) 1.98 (0.24)
H2OGBM_ElasticNet Jaccard 0.61 (0.19) 0.61 (0.20) 0.61 (0.20)
H2OGBM_ElasticNet Recovery 0.83 (0.07) 0.84 (0.08) 0.86 (0.09)
It can be seen that the hybrid models present an effective balance between predictive performance and robustness in variable selection, as in the CatBoost_Ridge model, which achieved the lowest root-mean-square error with $RMSE=1.89$ for $n=1000$, while CatBoost_Lasso stood out with the lowest RMSE for $n=200$ ($2.37$). The CatBoost_ElasticNet model presented the lowest RMSE with $n=500$ ($2.11$), in addition to achieving one of the best recovery rates ($0.86$).
Evaluating the Jaccard index, the best results were achieved by models associated with Lasso and Ridge, with CatBoost_Lasso standing out for $n=200$ and $n=500$, with $J=0.62$ and $J=0.61$, respectively. In the scenario with $n=1000$, the highest Jaccard was obtained by CatBoost_Ridge ($0.62$).
The recovery metric presented the highest values ($0.86$) across multiple models, such as CatBoost_ElasticNet, CatBoost_Ridge, and H2OGBM_ElasticNet, reflecting the increasing stability of variable selection with increasing sample size.
These results reinforce the value of hybrid models in practice, as they combine parsimony in selection with competitive performance in prediction, especially in scenarios with larger dimensions and larger samples.
3.4 Comparative Analysis and Implications
Black-box models excel in predictive accuracy (lower RMSE) but fail in interpretable selection, while pure regularized models offer good interpretability but present limitations under high spurious dimensionality, with high RMSEs and moderate Jaccard. Hybrid models occupy an intermediate position, combining good predictive performance and selection stability, especially as the sample size increases. Figure  1 visually demonstrates this intermediate position between accuracy and consistency.
Figure 1: Trade-off between predictive accuracy (RMSE) and selection consistency (Jaccard index) for the simulated models. Each point represents the average performance across 10 simulations for a model class. Error bars indicate one standard deviation.
The figure 1 illustrates the trade-off between accuracy and selection consistency among model classes. The black-box models obtained the lowest RMSEs, but with wider Jaccard indices compared to the others, demonstrating less stability. The regularized models presented the opposite pattern, still with high RMSE values and low Jaccard indices. Hybrid models occupy an intermediate position, balancing predictive accuracy with a wide range of simulations and selection stability.
In summary, hybrid models emerge as the most balanced solution for regression scenarios with high spurious dimensionality. They combine generalization, selection stability, and competitive performance. These findings are particularly relevant for applications in applied sciences and business environments, where accuracy and interpretability must go hand in hand.
4  Conclusion
The simulation results in different scenarios indicate that pure black-box models remain the most effective in terms of prediction error, but fail to promote parsimony. On the other hand, pure regularized models exhibit good stability in recovering relevant variables, but show limitations in predictive accuracy, with systematically higher RMSE values and only moderate Jaccard indices.
The hybrid models demonstrated remarkable performance by combining predictive accuracy with parsimony in variable selection. The combinations CatBoost_Lasso, CatBoost_ElasticNet, and H2OGBM_Ridge stood out, achieving low prediction errors while consistently identifying truly relevant variables. It was observed that, as the sample size (n) increases, there is a clear improvement in the retrieval and Jaccard metrics—indicating greater stability and accuracy in support reconstruction, even under high noise conditions.
Our findings reinforce the potential of hybrid models in practical applications, especially in high-dimensional contexts where the interpretability of results is as important as performance. By combining the selectivity of regularized methods with the flexibility of nonlinear algorithms, these approaches become robust and reliable solutions for real-world problems.
A possible future advancement is the creation of continuous and differentiable penalties integrated directly into the loss function of black-box models, such as CatBoost or XGBoost, which have shown good performance associated with the penalty, eliminating the need for external filtering steps. This would allow the development of truly hybrid models, with selection and prediction occurring in a unified manner.